# DreamTalk: When Emotional Talking Head Generation Meets Diffusion Probabilistic Models

### v2

@@ I Introduction
Existing methods still struggle to 1) consistently produce high-quality results across diverse speaking styles and 2) conveniently specify desired speaking styles. Existing methods [3, 5, 2, 6] are mainly based on GANs [7]. GANs’ inherent issues, such as mode collapse and unstable training, impair their performance across diverse speaking styles. Although these methods achieve satisfactory results for a limited range of speaking styles, they struggle with more diverse styles, especially ones unseen during training, often resulting in diminished emotional intensity, inaccurate lip motion, or sudden facial distortions [2]. Another issue is that to specify speaking styles, previous methods often rely on extra references, such as videos [3, 5, 1] or texts [8, 9]. Their acquisition requires extra manual effort and hence is inconvenient.
@@ II Related Work
Audio-Driven Talking Head Generation. Audio-driven methods [22, 23, 24, 25, 26, 27] fall into two main categories: person-specific and person-agnostic. Person-specific approaches [28, 29, 30, 31, 32] are constrained to generating videos for speakers seen during training. Many of these [33, 34, 35, 36, 37, 30, 38, 39, 40] first craft 3D facial animations, later converting them into realistic videos. Recent advancements [41, 42, 43, 44, 43, 45] have employed neural radiance fields for modeling, yielding high-fidelity, realistic videos. Conversely, person-agnostic methods [46, 47, 48, 49] target generating videos for unseen speakers. Early methods prioritized lip synchronization [50, 51, 52, 53, 49, 54]. Later works shifted focus to natural facial expressions [19, 55] and head poses [56, 57, 58, 59, 60]. FROND [61] introduces a fine-grained motion model that captures local facial movement keypoints and embeds overall motion context to predict audio-driven facial movements and achieve smooth temporal transitions. However, this method fails to generate emotional expressions during speech, thereby affecting the video’s realism.
Emotional Talking Head Generation. Early methods [9, 6, 62, 63, 64, 31, 30, 65, 66] model expressions in discrete emotions. To model fine-grained emotions, recent methods [5, 3, 2, 67] use an expression reference video and transfer the expressions from that video to the generated one. However, these GAN-based methods cannot consistently achieve high performance across diverse emotions. Our work addresses these issues by using diffusion models.
@@
Diffusion Models. Diffusion models [11, 10] have demonstrated strong performance across multiple vision tasks [12, 14, 71, 72, 73, 74], including text-to-image generation [75], human motion generation [17], and video generation [15, 76, 77, 78, 79]. Most diffusion-based methods for talking head generation [18, 80, 81, 82, 83, 84, 19, 85, 86, 87, 88], including EMO [21], AniPortrait [85] and Hallo [89], mainly generate talking heads with neutral emotions and lacks emotional controllability. Besides, the inference of EMO is slow. VASA [20] can only generate a limited number of emotions, lacking diverse, fine-grained speaking styles. In this work, we aim to harness diffusion models for generating and controlling diverse, fine-grained speaking styles in talking heads, presenting a more intricate challenge.
@@ V Limitations
DreamTalk may change the speaker’s identity when the identity in the reference video highly differs from that in the portrait. The reason is that 3DMM expression parameters leak identity information. This leakage leads to the generated identity becoming somewhat similar to the identity in the reference. The issue can be mitigated by adopting 3DMM which decouples expression and identity information more effectively.
The style predictor sometimes struggles with accurately identifying emotions in low-emotion-intensity audio clips from the MEAD dataset. The reason is that in some MEAD videos, the audio does not correspond with the expressed emotions. To enhance prediction accuracy, it is beneficial to employ a dataset where the audio closely aligns with the expressed emotions. Another solution is to incorporate textual information from audio during prediction, a strategy commonly employed in speech emotion recognition [105, 106].

### v3

@@ I Introduction
Existing methods still struggle to 1) consistently produce high-quality results across diverse speaking styles and 2) conveniently specify desired speaking styles. Existing methods [4, 6, 3, 7] are mainly based on GANs [8]. GANs’ inherent issues, such as mode collapse and unstable training, impair their performance across diverse speaking styles. Although these methods achieve satisfactory results for a limited range of speaking styles, they struggle with more diverse styles, especially ones unseen during training, often resulting in diminished emotional intensity, inaccurate lip motion, or sudden facial distortions [3]. Another issue is that to specify speaking styles, previous methods often rely on extra references, such as videos [4, 6, 2] or texts [9, 10, 11]. Their acquisition requires extra manual effort and hence is inconvenient.
@@ II Related Work
Audio-Driven Talking Head Generation. Audio-driven methods [24, 25, 26, 27, 28] fall into two main categories: person-specific and person-agnostic. Person-specific approaches [29, 30, 31, 32, 33] are constrained to generating videos for speakers seen during training. Many of these [34, 35, 36, 37, 31, 38, 39, 40] first craft 3D facial animations, later converting them into realistic videos. Recent advancements [41, 42, 43, 44, 43, 45] have employed neural radiance fields for modeling, yielding high-fidelity, realistic videos. Conversely, person-agnostic methods [46, 47, 48, 49] target generating videos for unseen speakers. Early methods prioritized lip synchronization [50, 51, 52, 53, 49, 54]. Later works shifted focus to natural facial expressions [21, 55] and head poses [56, 57, 58, 59, 60]. FROND [61] introduces a fine-grained motion model that captures local facial movement keypoints and embeds overall motion context to predict audio-driven facial movements and achieve smooth temporal transitions. However, this method fails to generate emotional expressions during speech, thereby affecting the video’s realism.
Emotional Talking Head Generation. Early methods [10, 7, 62, 63, 64, 32, 31, 65, 66] model expressions in discrete emotions. To model fine-grained emotions, recent methods [6, 4, 3, 1, 67] use an expression reference video and transfer the expressions from that video to the generated one. However, these GAN-based methods cannot consistently achieve high performance across diverse emotions. Our work addresses these issues by using diffusion models.
@@
Diffusion Models. Diffusion models [13, 12] have demonstrated strong performance across multiple vision tasks [14, 16, 71, 72, 73, 74], including text-to-image generation [75], human motion generation [19], and video generation [17, 76, 77, 78, 79]. Most diffusion-based methods for talking head generation [20, 80, 81, 82, 83, 84, 21, 85, 86, 87, 88, 89], including EMO [23], AniPortrait [85] and Hallo [90], mainly generate talking heads with neutral emotions and lacks emotional controllability. Besides, the inference of EMO is slow. VASA [22] can only generate a limited number of emotions, lacking diverse, fine-grained speaking styles. In this work, we aim to harness diffusion models for generating and controlling diverse, fine-grained speaking styles in talking heads, presenting a more intricate challenge.
@@ V Limitations
DreamTalk may change the speaker’s identity when the identity in the reference video highly differs from that in the portrait. The reason is that 3DMM expression parameters leak identity information. This leakage leads to the generated identity becoming somewhat similar to the identity in the reference. The issue can be mitigated by adopting 3DMM which decouples expression and identity more effectively.
The style predictor sometimes struggles with accurately identifying emotions in low-emotion-intensity audio clips from the MEAD dataset. The reason is that in some MEAD videos, the audio does not correspond with the expressed emotions. To enhance prediction accuracy, it is beneficial to employ a dataset where the audio closely aligns with the expressed emotions. Another solution is to incorporate textual information from audio during prediction, a strategy commonly employed in speech emotion recognition [106].